Problem 3-16 Comprehensive Problem [LO3-1, LO3-2, LO3-4]

Gold Nest Company of Guandong, China, is a family-owned enterprise that makes birdcages for the South China market. The company sells its birdcages through an extensive network of street vendors who receive commissions on their sales.

The company uses a job-order costing system in which overhead is applied to jobs on the basis of direct labor cost. Its predetermined overhead rate is based on a cost formula that estimated $85,500 of manufacturing overhead for an estimated activity level of $45,000 direct labor dollars. At the beginning of the year, the inventory balances were as follows:

Raw materials $ 10,600
Work in process $

4,800

Finished goods $ 8,400

During the year, the following transactions were completed:

  1. Raw materials purchased on account, $ 161,000.
  2. Raw materials used in production, $142,000 (materials costing $121,000 were charged directly to jobs; the remaining materials were indirect).
  3. Costs for employee services were incurred as follows:
Direct labor $ 167,000
Indirect labor $ 240,000
Sales commissions $ 28,000
Administrative salaries $

42,000

  1. Rent for the year was $18,000 ($13,900 of this amount related to factory operations, and the remainder related to selling and administrative activities).
  2. Utility costs incurred in the factory, $12,000.
  3. Advertising costs incurred, $13,000.
  4. Depreciation recorded on equipment, $20,000. ($17,000 of this amount related to equipment used in factory operations; the remaining $3,000 related to equipment used in selling and administrative activities.)
  5. Record the manufacturing overhead cost applied to jobs.
  6. Goods that had cost $226,000 to manufacture according to their job cost sheets were completed.
  7. Sales for the year (all paid in cash) totaled $508,000. The total cost to manufacture these goods according to their job cost sheets was $215,000.

Required:

1. Prepare journal entries to record the transactions for the year.

2. Prepare T-accounts for each inventory account, Manufacturing Overhead, and Cost of Goods Sold. Post relevant data from your journal entries to these T-accounts (don’t forget to enter the beginning balances in your inventory accounts).

3A. Is Manufacturing Overhead underapplied or overapplied for the year?

3B. Prepare a journal entry to close any balance in the Manufacturing Overhead account to Cost of Goods Sold.

4. Prepare an income statement for the year. All of the information needed for the income statement is available in the journal entries and T-accounts you have prepared.
Answer #1

journal entries

S.no account title particulars debit ($) credit ($)
1 RAW material inventory 161000
Cash 161000
2 work in process 121000
Manufacturing overhead 21000
Raw material inventory 142000
3 work in process 167000
Manufacturing overhead 240000
Sales commission 28000
Administrative salaries 42000
Cash 477000
4 manufacturing overhead 13900
Rent expenses 4100
Cash 18000
5 manufacturing overhead 12000
Cash 12000
6 advertising expenses 13000
Cash 13000
7 manufacturing overhead 17000
Depreciation expenses 3000
Accumulated depreciation 20000
8 work in process(note) 317300
Manufacturing overhead 317300
9 finished goods 226000
Work in process 226000
10 cash 508000
Sales 508000
11 cost of goods sold 215000
Finished goods 215000

   Note

   Pre determine overhead rate=85500/45000=1.9 per direct labour

   =167000×1.9=3173000

2). T accounts

RAW MATERIAL INVENTORY

Beginning balance 10600 work in process 121000
Cash 161000 manufacturing overhead 21000
Balance 29600

  work in process

Beginning balance 4800 finished goods 226000
RAW MATERIAL INVENTORY 121000
Cash 167000
Manufacturing overhead 317300 balance 384100

manufacturing overhead

Raw material inventory 21000 work in process 317300
Cash 240000
Cash 13900
Cash 12000
Cash 17000
Balance (overapplied) 13400

FINISHED GOODS

Beginning balance 8400 cost of goods sold 215000
Work in process 226000
Balance 19400

  cost of goods sold

Finished goods 215000 manufacturing overhead (overapplied) 13400
Balance 201600

3A MANUFACTURING OVERHEAD OVERAPPLIED=. ACTUAL OVERHEAD-APPILED OVERHEAD

= 303900 - 317300= 13400

3B

Manufacturing overhead 13400
Cost of goods sold 13400

4. INCOME STATEMENT

     

Particular amount
Sales 508000
Less cost of goods sold (201600)

Gross profit

306400

Less expenses

Sales commission (28000)
Administrative salaries (42000)
Rent (4100)
Advertising (3000)
Depreciation (13000)

  ​​​​​

operating income 216300

. ALL THE BEST

PLEASE DO SUPPORT US